The loss due to an earthquake in a commercial building is modelled by a random variable X with density function f(x) =0.005(20−x), for 0 < x < 20 0, elsewhere Given that the fire loss exceeds 10, what is the probability that it exceeds 18
1
Expert's answer
2020-05-14T17:51:13-0400

"P\{X>18|X>10\}=\frac{P\{X>18, X>10\}}{P\{X>10\}}\\
P\{X>18, X>10\}=P\{X>18\}\\
P\{X>18|X>10\}=\frac{P\{X>18\}}{P\{X>10\}}\\
F(x)=\int_{-\infty}^xf(t)dt.\\
f(x)\text{ is a pdf, } F(x)\text{ is a CDF}.\\
\text{Let } 0<x<20.\text{ We have:}\\
F(x)=\int_{0}^x 0.005(20-t)dt=-0.005\frac{x^2}{2}+0.1x.\\
F(x)=0, x<0.\\
F(x)=1, x>20.\\
P\{X>18\}=1-P\{X\leq 18\}=1-F(18)=\\
=1+0.005\frac{18^2}{2}-0.1\cdot 18=0.01.\\
P\{X>10\}=1-P\{X\leq 10\}=1-F(10)=\\
=1+0.005\frac{10^2}{2}-0.1\cdot 10=0.25.\\
P\{X>18|X>10\}=\frac{0.01}{0.25}=0.04\\
\text{Probability that the loss due to an earthquake}\\
\text{will exceed 18 equals 0.04 (given that the loss exceeds 10)}."